What is the mass of 764 milliliters of carbon disulfide, which has a density of 1.35 g/mL?

____ g
Answer #1

Answer

1031.4g

Explanation

Density = mass / volume

Density of CS2 = 1.35g/ml

given mass of CS2 = 764ml

1.35g/ml = mass/764ml

mass = 764ml × 1.35g/ml = 1031.4g